THz SPECTROSCOPY OF H_{2}D^{+}

\underline{F. MATSUSHIMA}, T. YONEZU, Y. MORIWAKI, and K. TAKAGI, Department of Physics, University of Toyama, Gofuku, Toyama 930-8555, Japan; T. AMANO, Department of Chemistry, University of Waterloo, Waterloo, Canada N2L 3G1.

The pure rotational transition frequencies of H_{2}D^{+}, 2_{12} ← 1_{11}, 2_{02} ← 1_{01}, 2_{11} ← 1_{10}, and 3_{13} ← 2_{12}, have been measured in the laboratory precisely by using a tunable far-infrared spectrometer. Among them, the 2_{12} ← 1_{11} line was recently detected in space toward Sgr B2 by Cernicharo et al.^{a}. Their identification was made based on a calculated line frequency estimated from the spectroscopic data of Amano and Hirao^{b}. It has been found that our measured frequency of this line, 2363242.82(69) MHz, is lower by about 20 MHz than the estimated value. All the available THz lines and known millimeter- and submillimeter-wave lines together with the combination differences derived from the infrared transitions are fitted to the Watson effective Hamiltonian. A set of improved molecular constants are obtained. —————————————- aJ.Cernicharo, E. Polehampton, and J.R. Goicoechea. Astrophys. J., 657, L21-L24 (2007). bT. Amano and T. Hirao, J. Mol. Spectrosc., 233, 7-14 (2005).